# Diffusion Models for Image Restoration and Enhancement: A Comprehensive Survey

### v2

@@ 6.3 Distortion Simulation and Estimation
Real-world/blind IR is a challenging yet significant task, with the goal of addressing the unknown and intricate degradations encountered in the real world. Unlike synthetic degradation, where the distortion is predefined and paired training samples are available, collecting the paired real-world distorted/clean pairs is non-trivial, thereby preventing the training with supervised learning. To address this limitation, unsupervised learning has been introduced to leverage unpaired real-world distorted/clean images. However, this learning paradigm usually yields an unsatisfied texture consistency between the restored image and the low-quality image. In contrast, distortion simulation serves as another efficient strategy to maintain supervised learning by simulating real-world degradations. Typically, RealESGRAN (Wang et al., 2021b) and SR3+ (Sahak et al., 2023) are the representative work to explore the hand-crafted second-order degradations for Real-world IR. Notwithstanding, the hand-crafted distortion simulation is hard to cover all the degradations in the real world. To mitigate this, some works are inspired by domain translation, and introduce the GAN/diffusion model to translate the synthetic distorted image to the real-world one or translate the real-world distorted image to the synthetic one. The former intends to simulate the real-world training pairs for supervised learning (Yang et al., 2023b), while the latter aims to directly utilize the IR network trained with synthetic images (Wei et al., 2023).
Another crucial challenge for real-world/blind IR stems from distortion estimation, which involves explicitly/implicitly identifying the distortion types or levels. In this work, we summarize the utilization of distortion estimation from two perspectives: 1) distortion adaptive learning and 2) solving the inverse problem in IR. From the first perspective, a notable example is kernel prediction for blind IR (Chung et al., 2023a), where the estimated kernel/representation is used to guide the adaptation of the pre-trained IR model to the unknown degradation. Inspired by this, if we can estimate the distortion type or degree in an explicit/implicit manner, we can achieve the unified IR framework based on distortion adaptive learning. For the second perspective, as stated in Sec. 3.4, lots of zero-shot diffusion model-based IR methods are based on the modelling of linear reverse problem. This poses the requirements for the identifying of degradation mode, which is necessary for the consistency constrain in diffusion model. Therefore, most of them are devoted into the synthetic distortions since the distortion mode in real world are hard to be identify. One distortion estimation technique is urgently developed to extend the zero-shot diffusion model-based IR to the real-world application.
@@ 6.4 Distortion Invariant Learning
Recently, we have witnessed the fast evolution of diffusion model-based IR for specific degradation. However, it inevitably suffers from unsatisfied robustness when applied to unseen distortion types and degrees. This raises a foundational question: How to achieve consistent image restoration across diverse distortion types and levels? To achieve this, we propose one direction termed distortion invariant learning (DIL) (Li et al., 2023c), aiming to enable the IR model to be generalized to unknown and diverse degradations. The principle of DIL is to learn the representation that is invariant under various degradation modes and preserve enough structure and textual information for reconstruction.
@@
Inspired by domain generalization (DG) (Zhou et al., 2022; Gulrajani and Lopez-Paz, 2020; Li et al., 2018a; Wang et al., 2022a), we can present some potential methods to achieve DIL for IR by regarding each distortion mode as one domain. In the DG field, there are three typical methods to learn the domain invariant feature, including domain alignment (Wang et al., 2021c; Li et al., 2018b; Shao et al., 2019), data augmentation (Xu et al., 2020; Mancini et al., 2020), and meta-learning (Wang et al., 2020a; Jia et al., 2020; Rahman et al., 2020; Albuquerque et al., 2019). In particular, domain alignment aims to align the representations of source and target domains through minimizing contrastive loss (Wang and Liu, 2021), Maximum Mean Discrepancy (MMD), or adversarial learning, etc. Data augmentation is exploited to extend the domain diversity and consistency, which enables the model to obtain the domain-invariant capability. Meta-learning aims to learn the domain invariant representation by aligning the gradient between different domains, which is from the optimization perspective. By regarding the distortion mode as one specific domain, we can obtain several strategies to achieve distortion invariant learning: 1) we can exploit the encoder-decoder architecture for IR and align the representations from different distorted images before the decoder. 2) the second strategy can learn the distortion invariant representation from distortion augmentation, i.e., simulating various distortions in the real world as much as possible. 3) optimize the empirical risk minimization in IR with meta-learning like (Li et al., 2023c).
@@
For diffusion model based image restoration, the model is usually composed of two components, noise predictor and condition module. Therefore, we can achieve distortion invariant learning from two perspectives: 1) learning the distortion invariant noise predictor and 2) distortion invariant condition. Obviously, once we implement the distortion invariant condition, we can leave the noise predictor invariant in supervised IR or exploit the pre-trained diffusion model in zero-shot IR. Based on this, some pioneering works attempt to redesign the condition module to achieve the distortion invariant condition, e.g., DifFace (Yue and Loy, 2024) and DR2 (Wang et al., 2023i). Notably, the distortion invariant condition also relies on distortion invariant learning for better conditions, which still entails lots of effort for evolution in future work.
@@ 6.5 Framework Design
As the foundation of image restoration, how to design an effective and powerful IR framework is an ongoing and significant question. We can notice that the most recent diffusion model-based IR methods (Saharia et al., 2022b; Wang et al., 2023g; Kawar et al., 2022a; Chung et al., 2023b; Saharia et al., 2022a; Song et al., 2023; Fei et al., 2023; Li et al., 2022b) are designed based on the U-Net architecture from DDPM (Ho et al., 2020), and pursue better frameworks from three perspectives, i.e., the condition strategy (Chung et al., 2023b; Kawar et al., 2022a; Lugmayr et al., 2022; Zhang et al., 2022a; Luo et al., 2023b), generation space(Rombach et al., 2022; Luo et al., 2023c; Xia et al., 2023), noise predictor (Xia et al., 2023), respectively. The condition of the diffusion model in IR aims to introduce the structure and textual information from low-quality images. In the early work, SR3 directly select low-quality images as a condition with the concatenation. To improve the condition, some works (Niu et al., 2023; Guo et al., 2023a; Gao et al., 2023b; Jin et al., 2022; Shang et al., 2023a) improve the condition by designing pre-processing networks, such as feature extractors and pre-trained restoration networks. From the generation space, the framework is usually designed from four spaces, including image space, residual space, latent space, and frequency space. Among them, pixel-wise space can preserve more spatial structure and textual information, which can generate high-quality images (Niu et al., 2023; Shang et al., 2023a; Gao et al., 2023b) or residuals (Li et al., 2022b; Shang et al., 2023a; Liu et al., 2024b; Yue et al., 2023; Liu et al., 2023a), while owning higher computational cost and parameters. In contrast, latent space generation requires fewer computational costs. However, one well-designed encoder and decoder are crucial for latent space generation to make the trade-off between efficiency and fidelity. Frequency space has been widely applied in diffusion model-based image rest-oration (Moser et al., 2023a; Xu et al., 2023; Mirza et al., 2023; Wang et al., 2023b; Guan et al., 2023; Luo et al., 2023a), including wavelet transform, Fourier transform, etc. Compared with image-wise space, frequency space is more good at capturing the global contextual information, where the low-frequency refers to the structural information and the high-frequency represents the texture and style information. These methods can bring more detailed and high-quality image restoration. Following the DDPM (Ho et al., 2020), in most works, the noise predictor is based on the U-Net architecture. For supervised diffusion model-based IR, the modification of the noise predictor is usually achieved by increasing the number of residual blocks in the U-Net or adjusting the channel multipliers at different resolutions, such as SR3. Few works are devoted to investigating how to design brand-new architectures based on transformers for the noise predictor in diffusion model-based IR. Furthermore, how to design the unified and foundational architecture like the painter (Wang et al., 2023f) for diffusion model-based IR tasks is urgently required to be explored.
@@ 6.6 All-in-one Restoration
All-in-one image restoration aims to simultaneously address multiple diverse degradations within a single IR model, which has attracted lots of attention in recent years. The key challenge lies in enabling diffusion-based IR models to adapt to various degradations while effectively leveraging shared restoration knowledge across different degradation types. The typical work is DA-CLIP (Luo et al., 2024), which introduces the degradation-aware CLIP to extract the degradation-aware embedding as the condition for diffusion models. In contrast, MPerceiver (Ai et al., 2024) extracts the holistic representation and multi-scale detail representation with the textual and visual branches of CLIP as the condition. Diff-Restore (Zhang et al., 2024) also leverages the CLIP model to extract the degradation and semantic embedding but adopts a distinct conditioning strategy. For unknown hybrid degradation, AutoDIR (Jiang et al., 2024b) introduce the quality assessment model to identify which degradation should be removed and utilizes multiple steps for restoration. UIR-LoRA (Hu et al., 2022) leverages low-rank adaptation (LoRA) to encode task embeddings efficiently. Then it introduces a dynamic weighting strategy based on CLIP similarity between image features and predefined distortion text embeddings. In contrast, GenDeg (Rajagopalan et al., 2024) utilizes diffusion models to generate diverse distortions from clean images, expanding training data for all-in-one diffusion-based IR models, which further improves the performance of multiple IR tasks. Although the above works have achieved impressive progress in diffusion-based all-in-one image restoration, a new all-in-one diffusion paradigm needs to be developed based on efficient architecture and new multi-task learning paradigm like Mixtures-of-Experts (MoE).
@@
Recent studies have proposed agent-based image restoration frameworks to enhance system intelligence, flexibility, and human interactivity. Leveraging the powerful understanding and reasoning capabilities of large language models (LLMs) and large multimodal models (LMMs), these approaches introduce language-driven agents into the restoration pipeline. For instance, AgenticIR (Zhu et al., 2024), hybridAgents Li et al. (2025) and RestoreAgent (Chen et al., 2024a) employ frozen LLMs or vision-language models to analyze input degradations and dynamically select traditional restoration tools (e.g., SwinIR (Liang et al., 2021), FBCNN (Jiang et al., 2021b), Restormer (Zamir et al., 2022)). Building upon this foundation, recent works such as MAIR (Jiang et al., 2025) and 4K-Agent (Yusheng et al., 2025) integrate powerful diffusion-based models—such as DiffBIR (Lin et al., 2024), IR-SDE (Luo et al., 2023b), and DPS (Chung et al., 2023b)—into the agent’s toolset. In particular, 4K-Agent employs a frozen LLM/MLLM to perform multi-round reasoning and sequential tool invocation, treating diffusion models as central, not auxiliary, components. This design achieves superior performance, especially under compound, real-world degradations. These developments reveal a clear trajectory in image restoration: from static, task-specific models to intelligent agentic systems, and further to fully integrated multimodal frameworks that emphasize interpretability, generalization, and controllability. As such, the development of LLM/MLLM-based restoration age-nts is emerging as a highly promising direction in the field, where the diffusion-based IR models are expected to serve as core generative priors and cross-modality generation tools in this system.
@@
The integration of vision-language models (VLMs) into restoration workflows enables high-level semantic guidance for diffusion models. PASD (Yang et al., 2024b) and Upscale-A-Video (Zhou et al., 2024) utilize frozen VLMs like BLIP-2 (Li et al., 2023a) and LLaVA (Liu et al., 2024a) to extract text prompts from degraded images, which are then used to condition diffusion upsamplers. Diff-Restorer (Zhang et al., 2024) takes this further by injecting semantic and degradation embeddings into the diffusion process. However, since standard VLMs may be misled by visual noise, recent methods like SeeSR (Wu et al., 2024b) and SUPIR (Yu et al., 2024) adapt VLMs to be degradation-aware by retraining (e.g., RAM-based tagging in SeeSR) or introducing lightweight artifact removers. These methods demonstrate the semantic understanding with VLMs, and the VLM-based condition strategies will empower high-fidelity restoration, which is a potential direction for diffusion-based IR.
@@
Most existing image restoration (IR) models are optimized using L1, perceptual, or similar handcrafted losses, which may not accurately reflect human preferences or perceptual quality. To bridge this gap, recent works have begun to explore diffusion-based generation with human feedback. For instance, PrefPaint (Liu et al., 2024c) learns a reward model from human-anno-tated preference data and fine-tunes a pre-trained diffusion backbone via reinforcement learning. InstructRestore (Liu et al., 2025) enables region-specific, instruction-driven restoration by allowing users to specify both spatial regions and restoration strength using natural language. DSPO (Cai et al., 2025) introduces patch-level semantic preference optimization for real-world super-resolution, aligning local instance appearance with human feedback. In the future, the human-in-the-loop optimization is expected to open up an exciting direction, such as reinforcement learning with human feedback (RLHF), synthetic preference data generation via diffusion sampling, and fine-grained/region-aware diffusion-based IR restoration under natural-language control.

### v3

@@ 6.3 Distortion Simulation and Estimation
Real-world/blind IR is a challenging yet significant task, with the goal of addressing the unknown and intricate degradations encountered in the real world. Unlike synthetic degradation, where the distortion is predefined and paired training samples are available, collecting the paired real-world distorted/clean pairs is non-trivial, thereby preventing the training with supervised learning. To address this limitation, unsupervised learning has been introduced to leverage unpaired real-world distorted/clean images. However, this learning paradigm usually yields an unsatisfied texture consistency between the restored image and the low-quality image. In contrast, distortion simulation serves as another efficient strategy to maintain supervised learning by simulating real-world degradations. Typically, RealESGRAN (wang2021realesrgan, ) and SR3+ (SR3+, ) are the representative work to explore the hand-crafted second-order degradations for Real-world IR. Notwithstanding, the hand-crafted distortion simulation is hard to cover all the degradations in the real world. To mitigate this, some works are inspired by domain translation, and introduce the GAN/diffusion model to translate the synthetic distorted image to the real-world one or translate the real-world distorted image to the synthetic one. The former intends to simulate the real-world training pairs for supervised learning (yang2023synthesizing, ), while the latter aims to directly utilize the IR network trained with synthetic images (wei2023raindiffusion, ).
Another crucial challenge for real-world/blind IR stems from distortion estimation, which involves explicitly/implicitly identifying the distortion types or levels. In this work, we summarize the utilization of distortion estimation from two perspectives: 1) distortion adaptive learning and 2) solving the inverse problem in IR. From the first perspective, a notable example is kernel prediction for blind IR (blinddps, ), where the estimated kernel/representation is used to guide the adaptation of the pre-trained IR model to the unknown degradation. Inspired by this, if we can estimate the distortion type or degree in an explicit/implicit manner, we can achieve the unified IR framework based on distortion adaptive learning. For the second perspective, as stated in Sec. 3.4, lots of zero-shot diffusion model-based IR methods are based on the modelling of linear reverse problem. This poses the requirements for the identifying of degradation mode, which is necessary for the consistency constrain in diffusion model. Therefore, most of them are devoted into the synthetic distortions since the distortion mode in real world are hard to be identify. One distortion estimation technique is urgently developed to extend the zero-shot diffusion model-based IR to the real-world application.
@@ 6.4 Distortion Invariant Learning
Recently, we have witnessed the fast evolution of diffusion model-based IR for specific degradation. However, it inevitably suffers from unsatisfied robustness when applied to unseen distortion types and degrees. This raises a foundational question: How to achieve consistent image restoration across diverse distortion types and levels? To achieve this, we propose one direction termed distortion invariant learning (DIL) (li2023learningDIL, ), aiming to enable the IR model to be generalized to unknown and diverse degradations. The principle of DIL is to learn the representation that is invariant under various degradation modes and preserve enough structure and textual information for reconstruction.
@@
Inspired by domain generalization (DG) (zhou2022domain-dg1, ; gulrajani2020search-dg2, ; li2018learning-dg3, ; wang2022generalizing-dg4, ), we can present some potential methods to achieve DIL for IR by regarding each distortion mode as one domain. In the DG field, there are three typical methods to learn the domain invariant feature, including domain alignment (9412797-alignment1, ; Li_2018_ECCV-alignment3, ; Shao_2019_CVPR-alignment4, ), data augmentation (xu2020robust-augmentation4, ; mancini2020towards-augmentation1, ), and meta-learning (wang2020meta-meta1, ; jia2020single-meta2, ; rahman2020correlation-meta3, ; albuquerque2019generalizing-meta4, ). In particular, domain alignment aims to align the representations of source and target domains through minimizing contrastive loss (wang2021understanding-contrastive-loss, ), Maximum Mean Discrepancy (MMD), or adversarial learning, etc. Data augmentation is exploited to extend the domain diversity and consistency, which enables the model to obtain the domain-invariant capability. Meta-learning aims to learn the domain invariant representation by aligning the gradient between different domains, which is from the optimization perspective. By regarding the distortion mode as one specific domain, we can obtain several strategies to achieve distortion invariant learning: 1) we can exploit the encoder-decoder architecture for IR and align the representations from different distorted images before the decoder. 2) the second strategy can learn the distortion invariant representation from distortion augmentation, i.e., simulating various distortions in the real world as much as possible. 3) optimize the empirical risk minimization in IR with meta-learning like (li2023learningDIL, ).
@@
For diffusion model based image restoration, the model is usually composed of two components, noise predictor and condition module. Therefore, we can achieve distortion invariant learning from two perspectives: 1) learning the distortion invariant noise predictor and 2) distortion invariant condition. Obviously, once we implement the distortion invariant condition, we can leave the noise predictor invariant in supervised IR or exploit the pre-trained diffusion model in zero-shot IR. Based on this, some pioneering works attempt to redesign the condition module to achieve the distortion invariant condition, e.g., DifFace (yue2022difface, ) and DR2 (wang2023dr2, ). Notably, the distortion invariant condition also relies on distortion invariant learning for better conditions, which still entails lots of effort for evolution in future work.
@@ 6.5 Framework Design
As the foundation of image restoration, how to design an effective and powerful IR framework is an ongoing and significant question. We can notice that the most recent diffusion model-based IR methods (SR3, ; ddnm, ; DDRMDRM, ; dps, ; saharia2022palette, ; iigdm, ; GDP, ; li2022srdiff, ) are designed based on the U-Net architecture from DDPM (DDPM, ), and pursue better frameworks from three perspectives, i.e., the condition strategy (dps, ; DDRMDRM, ; repaint, ; sr3_swinfir-transformer, ; IR-SDE, ), generation space(LDM, ; luo2023refusion, ; xia2023diffir, ), noise predictor (xia2023diffir, ), respectively. The condition of the diffusion model in IR aims to introduce the structure and textual information from low-quality images. In the early work, SR3 directly select low-quality images as a condition with the concatenation. To improve the condition, some works (niu2023cdpmsr, ; guo2023shadowdiffusion, ; IDM, ; jin2022shadowdiffusion-driven, ; shang2023resdiff, ) improve the condition by designing pre-processing networks, such as feature extractors and pre-trained restoration networks. From the generation space, the framework is usually designed from four spaces, including image space, residual space, latent space, and frequency space. Among them, pixel-wise space can preserve more spatial structure and textual information, which can generate high-quality images (niu2023cdpmsr, ; shang2023resdiff, ; IDM, ) or residuals (li2022srdiff, ; shang2023resdiff, ; liu2023residual-RDDM, ; yue2023resshift, ; liu2023textdiff, ), while owning higher computational cost and parameters. In contrast, latent space generation requires fewer computational costs. However, one well-designed encoder and decoder are crucial for latent space generation to make the trade-off between efficiency and fidelity. Frequency space has been widely applied in diffusion model-based image rest-oration (moser2023waving, ; xu2023stage, ; mirza2023learning-fourier-frequency-FDB, ; wang2023frequency-dehazy, ; guan2023correlated-frequency, ; luo2023image-frequency-feng, ), including wavelet transform, Fourier transform, etc. Compared with image-wise space, frequency space is more good at capturing the global contextual information, where the low-frequency refers to the structural information and the high-frequency represents the texture and style information. These methods can bring more detailed and high-quality image restoration. Following the DDPM (DDPM, ), in most works, the noise predictor is based on the U-Net architecture. For supervised diffusion model-based IR, the modification of the noise predictor is usually achieved by increasing the number of residual blocks in the U-Net or adjusting the channel multipliers at different resolutions, such as SR3. Few works are devoted to investigating how to design brand-new architectures based on transformers for the noise predictor in diffusion model-based IR. Furthermore, how to design the unified and foundational architecture like the painter (wang2023images-painter, ) for diffusion model-based IR tasks is urgently required to be explored.
@@ 6.6 All-in-one Restoration
All-in-one image restoration aims to simultaneously address multiple diverse degradations within a single IR model, which has attracted lots of attention in recent years. The key challenge lies in enabling diffusion-based IR models to adapt to various degradations while effectively leveraging shared restoration knowledge across different degradation types. The typical work is DA-CLIP (luo2023controlling-DACLIP, ), which introduces the degradation-aware CLIP to extract the degradation-aware embedding as the condition for diffusion models. In contrast, MPerceiver (ai2023multimodal-mperceriver, ) extracts the holistic representation and multi-scale detail representation with the textual and visual branches of CLIP as the condition. Diff-Restore (zhang2024diffrestore, ) also leverages the CLIP model to extract the degradation and semantic embedding but adopts a distinct conditioning strategy. For unknown hybrid degradation, AutoDIR (jiang2023autodir, ) introduce the quality assessment model to identify which degradation should be removed and utilizes multiple steps for restoration. UIR-LoRA (hu2022lora, ) leverages low-rank adaptation (LoRA) to encode task embeddings efficiently. Then it introduces a dynamic weighting strategy based on CLIP similarity between image features and predefined distortion text embeddings. In contrast, GenDeg (rajagopalan2024gendeg, ) utilizes diffusion models to generate diverse distortions from clean images, expanding training data for all-in-one diffusion-based IR models, which further improves the performance of multiple IR tasks. Although the above works have achieved impressive progress in diffusion-based all-in-one image restoration, a new all-in-one diffusion paradigm needs to be developed based on efficient architecture and new multi-task learning paradigm like Mixtures-of-Experts (MoE).
@@
Recent studies have proposed agent-based image restoration frameworks to enhance system intelligence, flexibility, and human interactivity. Leveraging the powerful understanding and reasoning capabilities of large language models (LLMs) and large multimodal models (LMMs), these approaches introduce language-driven agents into the restoration pipeline. For instance, AgenticIR (zhu2024intelligentAgenticIR, ), hybridAgents li2025hybrid and RestoreAgent (chen2024restoreagent, ) employ frozen LLMs or vision-language models to analyze input degradations and dynamically select traditional restoration tools (e.g., SwinIR (liang2021swinir, ), FBCNN (jiang2021towards-fbcnn, ), Restormer (zamir2022restormer_deblur-transformer, )). Building upon this foundation, recent works such as MAIR (jiang2025multiMAIR, ) and 4K-Agent (Zuo4KAgent, ) integrate powerful diffusion-based models—such as DiffBIR (lin2023diffbir, ), IR-SDE (IR-SDE, ), and DPS (dps, )—into the agent’s toolset. In particular, 4K-Agent employs a frozen LLM/MLLM to perform multi-round reasoning and sequential tool invocation, treating diffusion models as central, not auxiliary, components. This design achieves superior performance, especially under compound, real-world degradations. These developments reveal a clear trajectory in image restoration: from static, task-specific models to intelligent agentic systems, and further to fully integrated multimodal frameworks that emphasize interpretability, generalization, and controllability. As such, the development of LLM/MLLM-based restoration age-nts is emerging as a highly promising direction in the field, where the diffusion-based IR models are expected to serve as core generative priors and cross-modality generation tools in this system.
@@
The integration of vision-language models (VLMs) into restoration workflows enables high-level semantic guidance for diffusion models. PASD (yang2023pixel-PASD, ) and Upscale-A-Video (zhou2024upscaleavideo, ) utilize frozen VLMs like BLIP-2 (li2023blip2, ) and LLaVA (liu2024visual-LLaVA, ) to extract text prompts from degraded images, which are then used to condition diffusion upsamplers. Diff-Restorer (zhang2024diffrestore, ) takes this further by injecting semantic and degradation embeddings into the diffusion process. However, since standard VLMs may be misled by visual noise, recent methods like SeeSR (wu2024seesr, ) and SUPIR (yu2024scaling-SUPIR, ) adapt VLMs to be degradation-aware by retraining (e.g., RAM-based tagging in SeeSR) or introducing lightweight artifact removers. These methods demonstrate the semantic understanding with VLMs, and the VLM-based condition strategies will empower high-fidelity restoration, which is a potential direction for diffusion-based IR.
@@
Most existing image restoration (IR) models are optimized using L1, perceptual, or similar handcrafted losses, which may not accurately reflect human preferences or perceptual quality. To bridge this gap, recent works have begun to explore diffusion-based generation with human feedback. For instance, PrefPaint (liu2024prefpaint, ) learns a reward model from human-anno-tated preference data and fine-tunes a pre-trained diffusion backbone via reinforcement learning. InstructRestore (liu2025instructrestore, ) enables region-specific, instruction-driven restoration by allowing users to specify both spatial regions and restoration strength using natural language. DSPO (cai2025dspo, ) introduces patch-level semantic preference optimization for real-world super-resolution, aligning local instance appearance with human feedback. In the future, the human-in-the-loop optimization is expected to open up an exciting direction, such as reinforcement learning with human feedback (RLHF), synthetic preference data generation via diffusion sampling, and fine-grained/region-aware diffusion-based IR restoration under natural-language control.